# Parallelized Planning-Acting for Efficient LLM-based Multi-Agent Systems

### v1

@@ Abstract
Recent advancements in Large Language Model (LLM)-based Multi-Agent Systems (MAS) have demonstrated remarkable potential for tackling complex decision-making tasks. However, existing frameworks inevitably rely on serialized execution paradigms, where agents must complete sequential LLM planning before taking action. This fundamental constraint severely limits real-time responsiveness and adaptation, which is crucial in dynamic environments with ever-changing scenarios. In this paper, we propose a novel parallelized planning-acting framework for LLM-based MAS, featuring a dual-thread architecture with interruptible execution to enable concurrent planning and acting. Specifically, our framework comprises two core threads: (1) a planning thread driven by a centralized memory system, maintaining synchronization of environmental states and agent communication to support dynamic decision-making; and (2) an acting thread equipped with a comprehensive skill library, enabling automated task execution through recursive decomposition. Extensive experiments on challenging Minecraft demonstrate the effectiveness of the proposed framework.
Parallelized Planning-Acting for Efficient LLM-based Multi-Agent Systems
@@
1 Introduction
Figure 1: An overview of the proposed multi-agent system
Multi-Agent Systems (MAS) have become a well-established paradigm for tackling complex decision-making problems Hong et al. (2023); Chen et al. (2024b); Dong et al. (2024), with early efforts primarily relying on reinforcement learning Busoniu et al. (2008); Yang and Wang (2021); Lowe et al. (2017) to enable multiple agents to cooperate or compete in dynamic environments. Despite the encouraging results achieved, these MAS frameworks faced limitations in handling complex real-world scenarios that require advanced communication, reasoning, and adaptability. The rapid advancement of Large Language Models (LLMs) DeepSeek-AI (2025, 2024); Brown et al. (2020); Achiam et al. (2023); Qwen et al. (2024); Dubey et al. (2024) has since revolutionized MAS by adding natural language understanding and generation capabilities, enabling agents to engage in more sophisticated collaboration. LLMs have significantly enhanced the flexibility and versatility of MAS, opening the door to more complex tasks and dynamic interactions in real-world applications Wu et al. (2024); D’Arcy et al. (2024); Chen et al. (2024c); AL et al. (2024).
Recent works have demonstrated the potential of LLM-based MAS in various domains. MetaGPT Hong et al. (2023) introduces an innovative meta-programming framework to enhance task decomposition and agent collaboration, AgentVerse Chen et al. (2024b) improves collaborative performance by orchestrating expert agents, and VillagerAgent Dong et al. (2024) tackles task dependencies in complex environments using DAG-based task decomposition. Despite these advancements, most current frameworks still rely on serialized execution, where planning and acting occur sequentially for each agent. This serialized nature creates a substantial bottleneck when handling dynamic information, particularly evident in dynamic settings like Minecraft, where the environment is constantly changing, even without the intervention of agents. Although Voyager Wang et al. (2023) pioneered LLMs-based agents in Minecraft, its approach, which pauses the game server during LLM interactions, effectively staticizes the dynamic environment, thus failing to fully address the real-time responsiveness needed in such settings. Subsequent works in Minecraft agents Liu et al. (2024); Zhao et al. (2024) have followed this approach, but they also inherit the same limitations.
Our analysis reveals three critical challenges in current LLM-based MAS for dynamic environments. First, inflexible action scheduling is prevalent, as many existing agent frameworks rely on serialized execution, requiring agents to wait for a language model response before proceeding with further actions. This rigidity complicates the handling of unexpected environmental changes; Second, limited replanning capabilities hinder agents’ performance, as they often execute actions to completion without interruption. This lack of adaptability prevents agents from effectively reconsidering or adjusting their plans in response to urgent and unforeseen events, diminishing their overall effectiveness. Lastly, memory sharing delays pose another issue, as memory updates in many multi-agent systems only occur after an action has been fully executed. This results in delayed observational data sharing, causing agents to operate based on outdated information, which in turn limits the team’s coordination and efficiency.
In this paper, we propose a parallelized planning-acting framework that introduces a dual-thread architecture with interruptible execution for efficient LLM-based MAS in dynamic environments, as shown in Fig. 1. Our architecture decouples LLM reasoning from action execution, enabling concurrent planning and acting. Moreover, the interruption mechanism enables agents to adjust their actions in real time based on environmental changes, thereby improving their adaptability. Specifically, our framework consists of two core threads: (1) A planning thread employing a centralized memory system to support efficient and timely information sharing among agents, minimizing memory sharing delays and ensuring agents operate with up-to-date information for better coordination and efficiency. (2) An acting thread utilizing a comprehensive skill library, enabling efficient task execution through a recursive task decomposition mechanism. Our core contributions are summarized as follows:
  • We propose a parallelized planning-acting framework that decouples planning and acting into parallelized dual threads with interruptible execution for efficient LLM-based MAS.
  • We develop a centralized memory system to support the planning thread, ensuring that agent decisions are always informed by the latest environmental changes and interactions.
  • Experimental results on Minecraft demonstrate a paradigm shift from serialized deliberation to parallelized interaction, yielding notable improvements in efficiency and coordination.
2 LLM-based Multi-Agent Framework
We propose a novel parallelized planning-acting framework based on LLMs, designed to enable collaboration and decision-making among agents in complex dynamic environments. Our framework introduces three key innovations: (1) A dual-thread architecture with an interruptible execution mechanism, enabling concurrent planning and acting, (2) A real-time updated centralized memory system supporting the planning thread, ensuring that agents’ decisions are informed by the latest environmental changes and team interactions, and (3) A Minecraft comprehensive skill library supporting the acting thread, automating task execution by proposing a recursive task decomposition mechanism. While the skill library represents an engineering contribution, it is important to note that our framework is generic and can be extended to other domains with minimal adaptation. This flexibility highlights the broader applicability of our approach, while Minecraft provides a rich and accessible testbed for evaluating its performance in dynamic environments.
2.1 Parallelized Planning-Acting Framework
Inspired by the human ability to think and act simultaneously, our framework features a dual-thread architecture in Fig. 2 that separates planning (supported by LLMs and the centralized memory system) and acting (executed by the comprehensive skill library). Let $\mathcal{A}=\{a_{1},a_{2},\dots,a_{n}\}$ denote the set of agents. Our framework operates as follows:
[…]
    $A_{i}^{t+1}=\text{LLM}(S,O_{i}^{t},C^{t},A_{i}^{t}).$ (1)
    The planned action $A_{i}^{t+1}$ is then written into an action buffer, which serves as the communication channel between the planning thread and the acting thread. This action buffer is implemented as a single-slot queue, allowing the planning thread to write and the acting thread to read. If the buffer is already occupied, the previous action is discarded to make space for the new one. This mechanism ensures that the planning thread always places the most up-to-date action (based on the latest observations and agent status) into the buffer, while the acting thread always retrieves the most current action, never operating outdated plans.
  • Acting Thread: The acting thread retrieves actions from the action buffer and executes them according to the following rule:
    $s_{i}=\begin{cases}A_{i}^{t+1}&\text{if }p(A_{i}^{t+1})>p(A_{i}^{t}),\\ A_{i}^{t}&\text{otherwise}.\end{cases}$ (2)
    Here, $s_{i}$ represents the skill execution of agent $a_{i}$, $A_{i}^{t}$ is the current action being executed, $A_{i}^{t+1}$ is the new action in the buffer, and $p$ is the priority of an action. If an interruption is triggered—i.e., when the LLM determines that the new action has a higher priority than the current action—the planning thread sends an interrupt signal to the acting thread. The acting thread immediately exits and restarts with the new action. If no interruption is triggered, the acting thread completes the current action $A_{i}^{t}$ before retrieving and executing the next action $A_{i}^{t+1}$ from the buffer.
@@ 2.1 Parallelized Planning-Acting Framework
$\Delta T\approx\sum_{k=1}^{n}\left(T_{\text{plan}}^{(k)}+T_{\text{act}}^{(k)}-\max(T_{\text{plan}}^{(k)},T_{\text{act}}^{(k)})\right).$ (5)
This analysis highlights two key advantages of our framework: 1) The initial planning latency $T_{\text{plan}}^{(1)}$ is effectively amortized over subsequent actions, and 2) The overlapping of planning and acting phases successfully conceals $T_{\text{plan}}$ when $T_{\text{act}}>T_{\text{plan}}$ (our proposed comprehensive skill library in Section 2.3 ensures this condition is well-maintained). Our experimental results in Section 3.4 demonstrate the overall efficiency advantage of the parallelized framework.
By decoupling planning and acting into parallelized threads, our framework strikes a balance between efficiency and flexibility, allowing agents to respond dynamically to unpredictable environments with enhanced adaptability and performance.
2.2 Centralized Memory System
Figure 2: An overview of the proposed parallelized planning-acting architecture
@@ 2.2 Centralized Memory System
  • Observation Records: Each agent’s observations are continuously updated in the centralized memory, reflecting the latest agent status and environmental state. These observations are associated with the respective agent, allowing the team to maintain a comprehensive and up-to-date view of the environment.
@@
  • Chat Logs: All team chat messages are stored in the centralized memory, with long-term retention to support historical analysis and decision making. During planning, agents can retrieve the most recent chat messages to incorporate team insights into their strategies. This ensures that decisions are informed by the collective knowledge of the team.
  • Action History: The centralized memory also records the actions taken by each agent, providing a detailed history of task execution. These logs are useful for behaviours refining and performance analysis.
@@
We implement two types of multi-agent communication (passive and active) ensuring that the centralized memory remains a dynamic resource for team coordination:
@@
  • Passive Communication: In the planning thread, after each planning cycle, the LLM generates a chat message based on the agent’s latest observations, which is then sent to the centralized memory’s chat logs. This ensures that passive communication, reflecting updated observations, can run concurrently with action execution. While an agent is performing actions, its observations are continuously updated and shared with the team, enabling real-time coordination based on the most current environmental information.
@@
  • Active Communication: In the acting thread, agents can actively choose to send chat messages by performing a chat action. This action allows the agent to share any information with teammates, updating the chat logs in real-time. This form of communication ensures that agents can respond dynamically and share critical information during action execution, facilitating efficient and up-to-date information exchange between agents.
2.3 Comprehensive Skill Library
@@
To enable seamless interaction between agents and the Minecraft environment, we develop a comprehensive skill library that encapsulates a wide range of in-game actions. The library provides high-level APIs for tasks such as resource collection, combat, exploration, and communication. For further technical details, please refer to Appendix B.
@@
A key feature of the comprehensive skill library is the implementation of a recursive task decomposition mechanism, which automates the completion of prerequisite tasks such as mining raw materials and crafting necessary tools. This automation ensures that agents can perform complex resource collection tasks with minimal manual intervention, enabling the automated collection of over 790 types of items in Minecraft, surpassing all existing methods  Wang et al. (2023); Zhu et al. (2023); Zhao et al. (2024); Liu et al. (2024).
@@
  • Edge set $E\subseteq V\times V$ encodes prerequisite task dependencies:
@@
    where $\text{pre}(v_{i})$ gives the prerequisite tasks for producing $t_{i}$.
  • Weight function $\phi:E\to\mathbb{Q}^{+}$ defines material conversion rates:
@@
    with $r_{ij}$ being the required quantity of $t_{j}$ per operation, and $n_{\text{out}}$ being the output quantity per operation.
@@
where $v_{j}^{(k)}$ denotes a task requiring $k$ units of $t_{j}$, and the base case $\Psi(v_{i})=\emptyset$ applies when $I(t_{i})\geq c_{i}$, with $I$ representing the current inventory state, which means the task is accomplished.
3 Experiments
3.1 Experiment Setup
We aim to address the following questions through further experimentation: (1) Can our comprehensive skill library enable agents to efficiently accomplish basic tasks within Minecraft’s open-world environment? (2) How well can our proposed multi-agent framework perform on highly challenging and complex tasks? (3) Can our proposed parallelized planning-acting framework better support multi-agent collaboration, competition, and human-agent interaction?

### v2

# Parallelized Planning-Acting for Multi-Agent LLM Systems in Minecraft

@@ Abstract.
Recent advancements in Large Language Model (LLM)-based Multi-Agent Systems (MAS) have demonstrated remarkable potential for tackling complex decision-making tasks. However, existing frameworks inevitably rely on serialized execution paradigms, where agents must complete sequential LLM planning before taking action. This fundamental constraint severely limits real-time responsiveness and adaptation, which is crucial in dynamic environments with ever-changing scenarios like Minecraft. In this paper, we propose a novel parallelized planning-acting framework for LLM-based MAS, featuring a dual-thread architecture with interruptible execution to enable concurrent planning and acting. Specifically, our framework comprises two core threads: (1) a planning thread driven by a centralized memory system, maintaining synchronization of environmental states and agent communication to support dynamic decision-making; and (2) an acting thread equipped with a comprehensive skill library, enabling automated task execution through recursive decomposition. Extensive experiments on Minecraft demonstrate the effectiveness of the proposed framework.
Key words and phrases:
Multi-Agent Systems; Large Language Models
doi:
EXAJ9853
11footnotetext: Corresponding author.
1. Introduction
Multi-Agent Systems (MAS) have become a well-established paradigm for tackling complex decision-making problems Hong et al. (2023); Chen et al. (2024c); Dong et al. (2024), with early efforts primarily relying on reinforcement learning Busoniu et al. (2008); Yang and Wang (2021); Lowe et al. (2017) to enable multiple agents to cooperate or compete in dynamic environments. Despite the encouraging results achieved, these MAS frameworks faced limitations in handling complex real-world scenarios that require advanced communication, reasoning, and adaptability. The rapid advancement of Large Language Models (LLMs) DeepSeek-AI (2025, 2024); Brown et al. (2020); Achiam et al. (2023); Qwen et al. (2024); Yang et al. (2025); Dubey et al. (2024); Kimi (2025) has since revolutionized MAS by adding natural language understanding and generation capabilities, enabling agents to engage in more sophisticated collaboration. LLMs have significantly enhanced the flexibility and versatility of MAS, opening the door to more complex tasks and dynamic interactions in real-world applications Wu et al. (2024); D’Arcy et al. (2024); Chen et al. (2024a); AL et al. (2024).
Recent works have demonstrated the potential of LLM-based MAS in various domains. AgentVerse Chen et al. (2024c) improves collaborative performance by orchestrating expert agents, and VillagerAgent Dong et al. (2024) tackles task dependencies in complex environments using DAG-based task decomposition. Despite these advancements, most current frameworks applied in dynamic environments still rely on serialized execution, where planning and acting occur sequentially for each agent. This serialized nature creates a substantial bottleneck when handling dynamic information, particularly evident in dynamic settings like Minecraft, a game that features a vast and diverse world with various terrains, resources and creatures. Such an environment serves as an ideal testbed for evaluating the capabilities of MAS in open-world scenarios due to its constant environmental changes and rich interaction possibilities. While Voyager Wang et al. (2023) pioneered LLM-based agents in Minecraft, it relies on pausing the game server during planning to staticize the environment for the agent in order to prevent interference from environmental changes. However, in real-world dynamic environments, especially in multi-agent systems, it is not feasible to halt the actions of other agents while one agent is planning. This limitation hinders real-time interaction and reduces the system’s adaptability to sudden environmental changes or incoming information from other agents during LLM invocations.
Our analysis reveals three critical challenges in current LLM-based MAS for dynamic environments. First, inflexible action scheduling is prevalent, as many existing agent frameworks rely on serialized execution, requiring agents to wait for a language model response before proceeding with further actions. This rigidity complicates the handling of unexpected environmental changes. Second, limited replanning capabilities hinder agents’ performance, as they often execute actions to completion without interruption. This lack of adaptability prevents agents from effectively reconsidering or adjusting their plans in response to urgent and unforeseen events, diminishing their overall effectiveness. Lastly, memory sharing delays pose another issue, as memory updates in many multi-agent systems only occur after an action has been fully executed. This results in delayed observational data sharing, causing agents to operate based on outdated information, which in turn limits the team’s coordination and efficiency.
In this paper, we propose a parallelized planning-acting framework that introduces a dual-thread architecture with interruptible execution for efficient LLM-based MAS in dynamic environments. Our architecture decouples LLM reasoning from action execution, enabling concurrent planning and acting. Moreover, the interruption mechanism enables agents to adjust their actions in real time based on environmental changes, thereby improving their adaptability. Specifically, our framework consists of two core threads: (1) A planning thread employing a centralized memory system to support efficient and timely information sharing among agents, minimizing memory sharing delays and ensuring agents operate with up-to-date information for better coordination and efficiency. (2) An acting thread utilizing a comprehensive skill library, enabling efficient task execution through a recursive task decomposition mechanism. Our core contributions are summarized as follows:
  • We propose a parallelized planning-acting framework that decouples planning and acting into dual threads with interruptible execution for efficient LLM-based MAS.
  • We develop a centralized memory system to support the planning thread, ensuring agent decisions are informed by the latest environmental changes and interactions.
  • Experimental results on Minecraft demonstrate a paradigm shift from serialized deliberation to parallelized interaction, yielding notable improvements in efficiency, coordination and adaptability in dynamic environments.
2. LLM-based Multi-Agent Framework
We propose a novel parallelized planning-acting framework for LLM-based multi-agent systems as shown in Fig. 1, designed to support real-time inter-agent collaboration in dynamic scenarios such as the open-world environment of Minecraft. Our framework introduces three key innovations: (1) A dual-thread architecture with an interruptible execution mechanism, enabling concurrent planning and acting, (2) A real-time updated centralized memory system supporting the planning thread, ensuring that agents’ decisions are informed by the latest environmental changes and team communications, and (3) A comprehensive skill library supporting the acting thread, automating task execution by proposing a recursive task decomposition mechanism.
Figure 1. An overview of our parallelized planning-acting framework. In the multi-agent system, each agent operates with independent planning and acting threads in parallel. Planning threads are supported by a centralized memory system for decision-making, while acting threads utilize a DAG-based comprehensive skill library for task execution. After each planning step, the latest planned action is stored in the action buffer, and the agent performs a communication. Whenever an action is completed or the interrupt mechanism is triggered, the agent will continue to execute the skill function in the action buffer.
2.1. Parallelized Planning-Acting Framework
Inspired by the human ability to think and act simultaneously, our framework adopts a dual-thread architecture (Fig. 1) that decouples planning (driven by LLMs and the centralized memory system) from acting (executed by a comprehensive skill library). Let $\mathcal{G}=\{g_{1},g_{2},\dots,g_{n}\}$ denote the set of agents. The planning and acting threads operate asynchronously and independently, communicating only through a shared action buffer.
  • Planning Thread: The planning thread continuously monitors the environment and generates new action proposals for agent $g_{i}$ based on the system prompt $S$, the agent’s current observation $O_{i}$, the latest team chat logs $C$, and its current action $A_{i}$. At any time, the LLM may propose a new action together with an interruption flag:
    $A_{i}^{\text{new}},flag_{intr}=\text{LLM}(S,O_{i},C,A_{i}).$ (1)
    This proposed action $A_{i}^{\text{new}}$ is then written into a shared action buffer, which acts as a communication channel between the planning and acting threads. The buffer is implemented as a single-slot queue: if it is already occupied, the previous action will be overwritten. This ensures that the buffer always holds the most up-to-date action recommendation from the planner, reflecting the latest context. The interruption mechanism is also completely controlled by the LLM: it may trigger a restart if the new action is judged more urgent, or if the current action is no longer meaningful. When $flag_{intr}=\text{True}$, the planner issues a restart signal to the acting thread, ensuring that ongoing execution will be terminated and replaced by the new plan.
  • Acting Thread: The acting thread is responsible for executing skills from the comprehensive skill library. Let $A_{i}^{\text{curr}}$ denote the action currently being executed, and $A_{i}^{\text{new}}$ the action in the buffer. Upon restart, the acting thread immediately aborts its ongoing skill execution and fetches the latest action from the buffer. Otherwise, it continues executing $A_{i}^{\text{curr}}$ until completion, while periodically checking for updates. This design makes the acting thread reactive to planner-issued interrupts, rather than making its own decisions about preemption.
This decoupled design allows flexible, interruptible execution: planners can revise intentions at high frequency, while actors respond dynamically by replacing or preempting ongoing actions. Compared to fixed scheduling, this architecture significantly enhances responsiveness and adaptability in dynamic environments. Algorithm 1 summarizes the procedure.
Algorithm 1 Parallelized Planning-Acting Framework
1:Initialize agent set $\mathcal{G}\leftarrow\{g_{1},\dots,g_{n}\}$
@@ 2.1. Parallelized Planning-Acting Framework
4:procedure PlanningThread($g_{i}$)
5:  while not $\text{TaskCompleted}()$ do
6:   O_i $\leftarrow$ GetObservation(g_i)
7:   C $\leftarrow$ M.GetChatLogs()
8:   (new_action, intr) $\leftarrow$ LLM(S, O_i, C, cur_action)
9:   if $\texttt{intr}=True$ then
10:     Restart(ActingThread($g_{i}$))
11:   end if
12:   $Q.\text{put}(\texttt{new\_action},\text{overwrite=True})$
13:  end while
14:end procedure
15:procedure ActingThread($g_{i}$)
16:  while not $\text{TaskCompleted}()$ do
17:   if $Q.\text{has\_new}()$ then
18:     new_action $\leftarrow Q.\text{get}()$
19:     cur_action $\leftarrow$ new_action
20:   end if
21:   if $\texttt{cur\_action}\neq\text{None}$ then
22:     ExecuteSkill(cur_action)
23:     $M.\text{UpdateExecution}(\texttt{cur\_action})$
24:   end if
25:  end while
26:end procedure
27:Start PlanningThread and ActingThread for each $g_{i}\in\mathcal{G}$
28:Wait until all threads terminate
@@
$\begin{split}\Delta T&\approx\sum_{k}^{n}\left(T_{\text{plan}}^{(k)}+T_{\text{act}}^{(k)}-\max\left(T_{\text{plan}}^{(k)},T_{\text{act}}^{(k)}\right)\right)\\ &=\sum_{k}^{n}T_{\text{plan}}^{(k)}\quad\text{if for all }k,\,T_{\text{act}}^{(k)}>T_{\text{plan}}^{(k)}\end{split}$ (4)
This analysis highlights that the overlapping of planning and acting phases successfully conceals $T_{\text{plan}}$ when $T_{\text{act}}>T_{\text{plan}}$. We propose the comprehensive skill library in Section 2.3 to ensure this condition is well-maintained. For instance, a complex long-duration skill that might take several minutes to execute can be recursively decomposed and automated, while LLM reasoning typically only requires a few seconds. It is important to emphasize that our primary goal is to enable real-time and dynamic interaction, while latency reduction is a secondary benefit. Fig. 2 vividly presents a compact timeline diagram that illustrates the overlap between planning and acting, along with a concrete interrupt example.
$2.$ Centralized Memory System
@@
Figure 2. Schematic diagram of the timeline for parallel operation of the planning-acting dual threads. Among them, plans 2-3 are deemed not urgent enough to trigger the interrupt mechanism and thus are not actually executed. After the action 1 corresponding to plan 1 is completed, the latest action 4 corresponding to plan 4 is executed directly. Plan 6 sends an interrupt signal to suspend action 4, and the execution of action 6 starts immediately.
@@
  • Observation Records: Each agent’s observations are continuously updated in a polling manner (e.g. update per second), reflecting the latest agent status and environmental state. These observations are associated with the respective agent, allowing the team to maintain a comprehensive and up-to-date view of the environment.
@@
  • Chat Logs: Team chat messages are always updated in real time, with long-term retention to support historical analysis and decision making. During planning, agents can retrieve the most recent messages to incorporate team insights into their strategies, ensuring team collaboration.
  • Action History: Actions taken by each agent are also recorded, providing a detailed history of task execution. During planning, agents need to make decisions based on their current action and determine whether to interrupt it.
@@
We implement two types of multi-agent communication, including passive communication and active communication, ensuring a dynamic and diverse resource for team coordination among agents:
@@
  • Passive Communication: In the planning thread, the LLM generates a chat message based on the agent’s latest observations after planning, which is then sent to the centralized memory’s chat logs. This ensures that updated observations can run concurrently with action execution. While an agent is performing actions, its observations are continuously updated and shared with the team, enabling real-time coordination based on the most current environmental information.
@@
  • Active Communication: In the acting thread, agents can actively choose to send chat messages by performing a chat action implemented by the comprehensive skill library. This allows the agent to share any information with teammates, updating the chat logs in real-time. This form of communication ensures that agents can respond dynamically and share critical information, rather than passively communicating after the current action is over.
By supporting both passive and active communication modes, this system effectively addresses the challenge of memory sharing latency, ensuring that agents always operate on the latest team knowledge. For concrete examples illustrating these concepts, please refer to Appendix.
2.3. Comprehensive Skill Library
@@
To enable seamless interaction between agents and the Minecraft environment, we develop a comprehensive skill library based on Mineflayer PrismarineJS (2023) that encapsulates a wide range of in-game actions. The library provides high-level APIs for tasks such as resource collection, combat, exploration, and communication. For further technical details, please refer to Appendix.
@@
Our comprehensive skill library implements a recursive task decomposition mechanism, which automates the completion of prerequisite tasks such as mining raw materials and crafting necessary tools. This automation ensures that agents can perform complex resource collection tasks with minimal manual intervention, enabling the automated collection of over 790 types of items in Minecraft, surpassing all existing methods Wang et al. (2023); Zhu et al. (2023); Zhao et al. (2024); Liu et al. (2024).
@@
  • Edge set $E\subseteq V\times V$ encodes task dependencies:
@@
    where $\text{pre}(v_{i})$ gives the prerequisite tasks of $t_{i}$.
  • Weight function $\phi:E\to\mathbb{Q}^{+}$ defines conversion rates:
@@
    with $r_{ij}$ being the required quantity of $t_{j}$ and $n_{\text{out}}$ being the output quantity per operation.
@@
where $v_{j}^{(k)}$ denotes a task requiring $k$ units of $t_{j}$, and the base case $\Psi(v_{i})=\emptyset$ applies when $I(t_{i})\geq c_{i}$, with II representing the current inventory state, which means the task is accomplished.
This recursive task decomposition mechanism effectively models task prerequisite relationships as a Directed Acyclic Graph (DAG), where complex tasks are dynamically decomposed into atomic subtasks through automated dependency resolution. Unlike conventional approaches that require explicit step-by-step navigation from initial states to target objectives, our implementation enables agents to directly invoke high-level skills while the system automatically handles the recursive resolution of all prerequisite conditions. This design allows our multi-agent system to remain efficient by offloading detailed task execution to the skill library while leveraging LLMs for strategic decision-making and dynamic prioritization.
Moreover, the recursive task decomposition mechanism can be generalized to other environments beyond Minecraft. It is intuitive that allowing an LLM to independently plan the task dependencies embedded within a DAG would not be as efficient as utilizing a fixed mechanism. Consider a Directed Acyclic Graph (DAG) $\mathcal{G}=(V,E)$ with $n=|V|$ nodes and $e=|E|$ edges, representing task dependencies. Let the shortest path length from the source node $v_{s}$ (a node with zero in-degree) to a target node $v_{t}$ (a node with zero out-degree) be denoted as $L(v_{s},v_{t})$. This path corresponds to the minimal sequence of prerequisite tasks that must be executed to achieve the final objective represented by $v_{t}$.
In traditional planning approaches where each task is sequentially inferred by an LLM, the number of required model calls is at least $L(v_{s},v_{t})+1$, i.e., one for each node along the critical path. In contrast, our mechanism resolves all prerequisite dependencies automatically through the DAG structure, requiring only a single LLM invocation for the high-level task $v_{t}$.
This significant reduction in LLM usage not only improves computational efficiency but also reduces potential error propagation across multiple reasoning steps. Therefore, our methodology offers a generalizable framework for integrating LLMs with structured task execution systems across diverse domains beyond Minecraft.
The comprehensive skill library also demonstrates strong scalability and can be readily extended to accommodate new updates. Newly introduced skills can be seamlessly integrated by simply updating the prerequisite dependency DAG that encodes task relationships without new interfaces development or extensive code modifications. This modular design ensures continued functionality with minimal maintenance effort, making the library both future-proof and adaptable to evolving domain requirements.
3. Experiments
3.1. Benchmark Task Design
Our experiments are designed to validate the framework’s capabilities, leveraging Minecraft as a testbed while focusing on general Multi-Agent System (MAS) competencies. Existing Minecraft agent methods suffer from a lack of a comprehensive skill library, which hinders their ability to learn complex strategies for tackling challenging tasks. Consequently, we have developed more demanding tasks built upon the existing evaluation paradigm, rather than limiting our scope to basic tasks solely for baseline comparisons. For example, defeating the Ender Dragon is widely recognized as Minecraft’s ultimate challenge, which remains unachievable with current methods to the best of our knowledge, serves as one of our evaluation tasks. Our benchmark comprises a diverse range of tasks and supports standardized evaluation for potential future research. Refer to the Appendix for more details.
  • Resource Collection Task: Evaluates the foundation of our framework through fundamental mining and crafting tasks in Minecraft. These tasks aim to validate the effectiveness of our comprehensive skill library by requiring agents to complete compound items with deep dependency chains and measure multi-agent coordination efficiency in distributed resource collection. we define a set of representative resource collection tasks, such as Diamond Armor serves as a foundation for challenging combat scenarios in the game.
  • Boss Combat Task: Assesses dynamic adaptation and strategical coordination through combat scenarios against powerful bosses. In Minecraft, there are three primary world dimensions: Overworld, Nether, and End. Each dimension hosts powerful boss monsters whose defeat is considered pinnacle challenges for players of Minecraft. Based on this, we define three representative combat tasks.
  • Adversarial PVP Task: Compares two frameworks directly by having them engage in direct combat scenarios. In this task, two teams of agents engage in battles with each other.
3.2. Experimental Setup